# Rapid and Power-Aware Learned Optimization for Modular Receive Beamforming

### v1

@@ V-B Unconstrained Modular Beamformers
We start the experimental study by considering unconstrained modular beamformers, as described in Subsection III-B. Our evaluation first focuses on small scale MIMO settings, after which we assess transferability to larger systems.
@@ V-B1 Small-Scale Hybrid Modular MIMO
For fair comparison, we set all optimization-based algorithms to have $J=10$ iterations, with the CNN having $10$ layers, and compare the sum-rate to that achieved by PGA+M with $J=500$ iterations. We consistently observe that in the considered scenarios with a limited number of iterations, our unfolded PGA outperforms all benchmarks, including the costly and greedy line search hyperparameter setting, and the highly parameterized CNN benchmark. The latter particularly struggles in Rayleigh channels, as CNNs tends to capture spatial relationships, and such relationships do not exist under Rayleigh fading. Similarly, the MO optimizer exhibits slow convergence due to the additional modular structure imposed in our setting of modular beamformers.
The results after $J$ iterations illustrated in Figs. 5-6 are also reported in Table I, along with latency of the various methods. The latter is evaluated for all algorithms on the same platform (Google Colab premium). We notice in Table I that the unfolded algorithm is $30-50$ times faster than optimizers that employ line search or run a large number of iterations, while achieving a comparable sum-rate. The CNN benchmark has the least latency here, owing to the efficient implementation and natural support of such architectures in platforms such as Google Colab. This efficiency persists across multiple scenarios, emphasizing our ability to provide reliable modular beamformers with minimal latency, that can be further enhanced by employing dedicated hardware accelerators.
@@ V-B2 Transferable Modular Structures Simulation Results
As explained in Subsection III-B3, one of the key advantages of our learn-to-optimize modular beamforming design is the transferability of learned parameters when scaling the system. Modular structures allow us to train the system with specific settings and then scale it up to a larger number of users or panels without needing to retrain the model.
@@
Scaling the number of panels: Next, we provide an example of scaling up the number of panels. In this example, we use $c=8$ as the scaling coefficient, and increase the number of panels from $P$ to $c\cdot P$, using $\{{\boldsymbol{\alpha}}_{j},{\boldsymbol{\beta}}_{j}\}=\{{\boldsymbol{I}}_{c}\otimes{\boldsymbol{\alpha}}_{j},{\boldsymbol{I}}_{c}\otimes{\boldsymbol{\beta}}_{j}\}$. To this end, we introduce Scenario 5: $(T=5,P=2,N=4,B=2.K=7)$ and Scenario 6: $(T=5\cdot 8,P=2\cdot 8,N=4,B=2,K=7)$; both obey Rayleigh fading.
@@
Fig. 10 illustrates the results of various benchmarks trained on Scenario 6, alongside the performance of the U-PGA+M algorithm that was trained on Scenario 5. These results highlight the flexibility of our design, demonstrating how it can be initially trained on a smaller network, perform straightforward calculations, and then be effectively scaled to much larger and more complex systems.
@@ V-C1 Sparse Modular Beamformers
Figure 12: Sum-rate vs. SNR, Scenario 2
@@ V-C2 Low-Resolution Modular Beamformers
The resulting sum-rate vs. iteration and vs. SNR for Scenario 1 are illustrated in Figs. 13 and 14, respectively. The results demonstrate that the proposed unfolded algorithm, which is particularly geared to accounting for low-resolution phase shifters, significantly outperforms other benchmarks, including both slow iterative methods that are not trained as machine learning models, and suffer more notable degradation due to quantization, as well as the heavily parameterized CNN. Specifically, while PGA + M struggles to overcome the quantization constraints using gradient steps, and tends to converge to a local maximum, the unfolded algorithm effectively learns to traverse the loss surface and projection constraints, achieving a better sum-rate. As shown in Fig. 14, the unfolded algorithm consistently performs better across various SNR values compared to all other benchmarks.
@@ VI Conclusions
For ease of presentation, we define the following auxiliary variables: ${\boldsymbol{G^{\dagger}}}\triangleq({\boldsymbol{G}}^{H}{\boldsymbol{G}})^{-1}{\boldsymbol{G}}^{H}$, ${\boldsymbol{P}}\triangleq{\boldsymbol{G}}{\boldsymbol{G^{\dagger}}}$, ${\boldsymbol{Z}}[b]\triangleq{\boldsymbol{P}}{\boldsymbol{H}}[b]{\boldsymbol{H}}^{H}[b]$, and ${\boldsymbol{Q}}[b]\triangleq{\boldsymbol{I}}_{k}+\frac{\rho_{s}}{\sigma_{w}^{2}}{\boldsymbol{Z}}[b]$. Using these symbols, the objective with respect to which the derivative is taken can be written as $R\left({\boldsymbol{G}};\{{\boldsymbol{H}}[b]\}\right)=\frac{1}{B}\sum_{b=1}^{B}\log\big{|}{\boldsymbol{Q}}[b]\big{|}$. Plugging-in auxiliary variables, we aim at proving Theorem 1 by showing that
@@
$dR=$ $\Big{(}{\boldsymbol{Q}}^{-1}\Big{)}^{H}\;\textbf{:}\;d{\boldsymbol{Q}}\stackrel{{\scriptstyle(a)}}{{=}}\Big{(}{\boldsymbol{Q}}^{-1}\Big{)}^{H}\;\textbf{:}\;\frac{\rho_{s}}{\sigma_{w}^{2}}d{\boldsymbol{Z}}$
=$\stackrel{{\scriptstyle(b)}}{{=}}$ $\Big{(}\frac{\rho_{s}}{\sigma_{w}^{2}}{\boldsymbol{Q}}^{-1}\Big{)}^{H}\;\textbf{:}\;d{\boldsymbol{Z}}\stackrel{{\scriptstyle(c)}}{{=}}\Big{(}\frac{\rho_{s}}{\sigma_{w}^{2}}{\boldsymbol{Q}}^{-1}\Big{)}^{H}\;\textbf{:}\;d{\boldsymbol{P}}\,{\boldsymbol{H}}{\boldsymbol{H}}^{H}$
=$\stackrel{{\scriptstyle(d)}}{{=}}$ $\Big{(}\frac{\rho_{s}}{\sigma_{w}^{2}}\,{\boldsymbol{H}}{\boldsymbol{H}}^{H}{\boldsymbol{Q}}^{-1}\Big{)}^{H}\;\textbf{:}\;({\boldsymbol{I}}-{\boldsymbol{P}})\,d{\boldsymbol{G}}\>{\boldsymbol{G}}^{\dagger}$ (-A.2)
@@
Next, we substitute ${\boldsymbol{P}}$, ${\boldsymbol{G}}^{\dagger}$, and ${\boldsymbol{G}}$ using chain rule:
@@
$=d{\boldsymbol{G}}\,{\boldsymbol{G}}^{\dagger}-{\boldsymbol{G}}({\boldsymbol{G}}^{H}{\boldsymbol{G}})^{-1}{\boldsymbol{G}}^{H}d{\boldsymbol{G}}\,{\boldsymbol{G}}^{\dagger}$
@@
In (-A), the term $d{\boldsymbol{G}}^{H}$ vanishes since ${\boldsymbol{G}}^{H}$ is independent of ${\boldsymbol{G}}$, and thus independent of ${\boldsymbol{W}}$. Once (-A.3) is achieved, the rest of the proof becomes straight forward, using the chain rule until (-A.1) is reached, i.e.,
@@
$=$ $\Big{(}\frac{\rho_{s}}{\sigma_{w}^{2}}\,{\boldsymbol{G}}^{\dagger}{\boldsymbol{H}}{\boldsymbol{H}}^{H}{\boldsymbol{Q}}^{-1}({\boldsymbol{I}}-{\boldsymbol{P}})\Big{)}^{H}\;\textbf{:}\;\,d{\boldsymbol{W}}{\boldsymbol{A}}$
$=$ $\Big{(}\frac{\rho_{s}}{\sigma_{w}^{2}}\,{\boldsymbol{A}}{\boldsymbol{G}}^{\dagger}{\boldsymbol{H}}{\boldsymbol{H}}^{H}{\boldsymbol{Q}}^{-1}({\boldsymbol{I}}-{\boldsymbol{P}})\Big{)}^{H}\;\textbf{:}\;\,d{\boldsymbol{W}},$ (-A.4)
which concludes that $\nabla_{\boldsymbol{W}}R({\boldsymbol{W}},{\boldsymbol{A}})=\big{(}\frac{\rho_{s}}{\sigma_{w}^{2}}{\boldsymbol{A}}{\boldsymbol{G^{\dagger}}}{\boldsymbol{H}}{\boldsymbol{H}}^{H}{\boldsymbol{Q}}^{-1}({\boldsymbol{I}}-{\boldsymbol{P}})\big{)}^{H}$, thus proving the theorem.

### v2

@@ V-B Unconstrained Modular Beamformers
We start the experimental study by considering unconstrained modular beamformers, as described in Subsection III-B. Our evaluation first focuses on small scale MIMO settings, after which we assess transferability to larger systems, and then we examine the performance of our robust algorithm in tackling CSI uncertainty, demonstrating its ability to maintain reliability and efficiency under imperfect channel knowledge.
@@ V-B1 Small-Scale Hybrid Modular MIMO
For fair comparison, we set all optimization-based algorithms to have $J=10$ iterations, with the CNN having $10$ layers, and compare the sum-rate to that achieved by PGA+M with $J=500$ iterations. We consistently observe that in the considered scenarios with a limited number of iterations, our unfolded PGA outperforms all benchmarks, including the costly and greedy line search hyperparameter setting, and the highly parameterized CNN benchmark. Specifically, the conventional PGA+M requires 6.7×6.7$\times$6.7 × (for Rayleigh fading) and 14.7×14.7$\times$14.7 × (for QuaDRiGa) more iterations to achieve the performance of our unfolded algorithm. The CNN benchmark particularly struggles in Rayleigh channels, as CNNs tends to capture spatial relationships, and such relationships do not exist under Rayleigh fading. Similarly, the MO optimizer exhibits slow convergence due to the additional modular structure imposed in our setting of modular beamformers.
[…]
@@ V-B2 Transferable Modular Structures Simulation Results
As explained in Subsection III-B4, one of the key advantages of our learn-to-optimize modular beamforming design is the transferability of learned parameters when scaling the system. Modular structures allow us to train the system with specific settings and then scale it up to a larger number of users or panels without needing to retrain the model.
@@
Scaling the number of panels: Next, we provide an example of scaling up the number of panels. In this example, we use $c=8$ as the scaling coefficient, and increase the number of panels from $P$ to $c\cdot P$, using $\{{\boldsymbol{\alpha}}_{j},{\boldsymbol{\beta}}_{j}\}=\{{\boldsymbol{I}}_{c}\otimes{\boldsymbol{\alpha}}_{j},{\boldsymbol{I}}_{c}\otimes{\boldsymbol{\beta}}_{j}\}$. To this end, we introduce Scenario 5: $(T=5,P=2,N=4,B=2,K=7)$ and Scenario 6: $(T=5\cdot 8,P=2\cdot 8,N=4,B=2,K=7)$; both obey Rayleigh fading.
@@
Fig. 10 illustrates the results of various benchmarks trained on Scenario 6, alongside the performance of the U-PGA+M algorithm that was trained on Scenario 5. While the gains of our unfolded algorithm (trained for Scenario 5) compared to the benchmarks that are all tuned for Scenario 6 are more modest here compared with having it trained for the same setup, the fact that it does not degrade performance, and is even applicable in such different setups without necessitating any form of retraining, is quite non-trivial. These results highlight the flexibility of our design, demonstrating how it can be initially trained on a smaller network, perform straightforward calculations, and then be effectively scaled to much larger and more complex systems.
V-B3 Robustness to CSI Estimation Error Simulation Results
As detailed in Subsection III-B3, the trainabiltiy of our unfolded optimizer can be leveraged to cope with CSI estimation errors caused by e.g., noise or limited pilots. To evaluate the robustness of our unfolded beamforming algorithm, we simulate the impact of CSI estimation errors in Scenario 2 with 00 dB SNR, and analyze its performance in terms of sum-rate as the variance of the estimation error increases.
The results, shown in Fig. 11, highlight the comparative robustness of the proposed unfolded PGA algorithm (U-PGA+M), which operates with solely 10 iterations. While performance degradation is observed for all methods as the estimation error variance ($\sigma_{E}^{2}$) increases, the unfolded PGA algorithm demonstrates a less steep descent compared to the conventional PGA+M approach, even when the latter is executed with 500 iterations. In particular, PGA+M with 10 iterations exhibits the most significant performance degradation, indicating that insufficient optimization steps amplify the sensitivity to CSI estimation errors. In contrast, the PGA+M method with 500 iterations reduces this sensitivity but still underperforms relative to the unfolded algorithm in scenarios with high $\sigma_{E}^{2}$. This indicates the ability of our unfolded approach in achieving robustness imperfect CSI via dedicated learnin. The unfolded algorithm’s performance at lower $\sigma_{E}^{2}$ remains competitive with conventional methods, confirming that the benefits of robustness do not come at the cost of baseline performance. Overall, these findings demonstrate that the proposed unfolded PGA framework not only provides robust performance under realistic channel conditions, but also offers a scalable solution for handling imperfect CSI in modular beamforming systems.
Figure 11: Sum-rate vs. CSI estimation error variance.
@@ V-C1 Sparse Modular Beamformers
Figure 13: Sum-rate vs. SNR, Scenario 2.
@@ V-C2 Low-Resolution Modular Beamformers
The resulting sum-rate vs. iteration and vs. SNR for Scenario 1 are illustrated in Figs. 14 and 15, respectively. The results demonstrate that the proposed unfolded algorithm, which is particularly geared to accounting for low-resolution phase shifters, significantly outperforms other benchmarks, including both slow iterative methods that are not trained as machine learning models, and suffer more notable degradation due to quantization, as well as the heavily parameterized CNN. Specifically, while PGA + M struggles to overcome the quantization constraints using gradient steps, and tends to converge to a local maximum even when allowed to run for $500$ iterations, the unfolded algorithm effectively learns to traverse the loss surface and projection constraints, achieving a better sum-rate. As shown in Fig. 15, the unfolded algorithm consistently performs better across various SNR values compared to all other benchmarks.
@@ VI Conclusions
For ease of presentation, we define the following auxiliary variables: ${\boldsymbol{G^{\dagger}}}\triangleq({\boldsymbol{G}}^{H}{\boldsymbol{G}})^{-1}{\boldsymbol{G}}^{H}$, ${\boldsymbol{P}}\triangleq{\boldsymbol{G}}{\boldsymbol{G^{\dagger}}}$, ${\boldsymbol{Z}}[b]\triangleq{\boldsymbol{P}}{\boldsymbol{H}}[b]{\boldsymbol{H}}^{H}[b]$, and ${\boldsymbol{Q}}[b]\triangleq{\boldsymbol{I}}_{k}+\frac{\rho_{s}}{\sigma_{w}^{2}}{\boldsymbol{Z}}[b]$. Using these symbols, the objective with respect to which the derivative is taken is written as $R\left({\boldsymbol{G}};\{{\boldsymbol{H}}[b]\}\right)=\frac{1}{B}\sum_{b=1}^{B}\log\big{|}{\boldsymbol{Q}}[b]\big{|}$. Substituting auxiliary variables, we aim at proving (9) by showing that
@@
$dR=$ $\Big{(}{\boldsymbol{Q}}^{-1}\Big{)}^{H}\;\textbf{:}\;d{\boldsymbol{Q}}\stackrel{{\scriptstyle(a)}}{{=}}\Big{(}\frac{\rho_{s}}{\sigma_{w}^{2}}{\boldsymbol{Q}}^{-1}\Big{)}^{H}\;\textbf{:}\;d{\boldsymbol{P}}\,{\boldsymbol{H}}{\boldsymbol{H}}^{H},$ (-A.2)
@@
Next, we substitute ${\boldsymbol{P}}$, ${\boldsymbol{G}}^{\dagger}$, and ${\boldsymbol{G}}$ using the chain rule:
@@
Once (-A.3) is achieved, the rest of the proof becomes straight forward, using the chain rule until (-A.1) is reached, i.e.,
@@
$=$ $\Big{(}\frac{\rho_{s}}{\sigma_{w}^{2}}\,{\boldsymbol{A}}{\boldsymbol{G}}^{\dagger}{\boldsymbol{H}}{\boldsymbol{H}}^{H}{\boldsymbol{Q}}^{-1}({\boldsymbol{I}}-{\boldsymbol{P}})\Big{)}^{H}\;\textbf{:}\;\,d{\boldsymbol{W}},$
which concludes that $\nabla_{\boldsymbol{W}}R({\boldsymbol{W}},{\boldsymbol{A}})=\big{(}\frac{\rho_{s}}{\sigma_{w}^{2}}{\boldsymbol{A}}{\boldsymbol{G^{\dagger}}}{\boldsymbol{H}}{\boldsymbol{H}}^{H}{\boldsymbol{Q}}^{-1}({\boldsymbol{I}}-{\boldsymbol{P}})\big{)}^{H}$, thus proving (9).